OpenStudy (anonymous):

Find the most general antiderivative of the function. F(x)= x^3^(1/4)+ x^8^(1/3)

OpenStudy (anonymous):

\[f(x)=\sqrt[4]{x^3}+\sqrt[3]{x^8}\] this is the problem written out with equation

OpenStudy (anonymous):

the way you have written it in your question is perfect, you can apply single integral calculus rules to it, (increase the exponent by +1 and divide by the new exponent)

OpenStudy (anonymous):

\[\large \int x^{\frac{3}{4}}dx+\int x^{\frac{8}{3}}dx \]

OpenStudy (anonymous):

ok I got it thanks.

OpenStudy (anonymous):

welcome